# An operator learning perspective on parameter-to-observable maps

### v1

@@ C.2. Proofs for end-to-end learning of general linear functionals
[…]
@@ C.2.1. Bounding the variance
The expression for the variance upper bound in Proposition C.5 is in precisely the same form as that found in [16, Equation (A.1b), p. 24]. We derive quantitative convergence rates for $V_{N}$ under the assumption of power law decay of the eigenvalues of the data and prior covariance operators. However, other types such as exponential decay [3] or convex eigenvalues [12] are also possible and interesting.
To prove the proposition, we require some preparatory results. First, notice that
$\operatorname{\mathbb{E}}\mathopen{}\mathclose{{}\left[\big{\lVert}(\Sigma^{\prime})^{1/2}A_{N}\Xi\big{\rVert}^{2}\,\big{|}\,U}\right]=\operatorname{\mathbb{E}}^{g\sim\mathcal{N}(0,\operatorname{Id}_{{\mathbb{R}^{N}}})}\big{\lVert}(\Sigma^{\prime})^{1/2}A_{N}g\big{\rVert}^{2}=\operatorname{tr}\big{(}(\Sigma^{\prime})^{1/2}A_{N}A_{N}^{*}(\Sigma^{\prime})^{1/2}\big{)}\,.$
Now using the fact that $\widehat{\Sigma}=S_{N}^{*}S_{N}/N$ from (14) and defining
@@
$A_{N}A_{N}^{*}=\frac{1}{N}\biggl{[}\Lambda^{1/2}\biggl{(}\widehat{\mathcal{C}}+\frac{\gamma^{2}}{N}\operatorname{Id}_{H}\biggr{)}^{-1}\widehat{\mathcal{C}}\biggl{(}\widehat{\mathcal{C}}+\frac{\gamma^{2}}{N}\operatorname{Id}_{H}\biggr{)}^{-1}\Lambda^{1/2}\biggr{]}\,.$
@@
The proof of Proposition C.5 also justifies the claim made in Remark 4.9. Indeed, the second term on the right hand side of the equality (33) is the posterior spread with respect to the weighted norm (18). Equation (73) upper bounds the posterior spread, which in turn upper bounds the variance (68b) in the bias–variance decomposition of (18). Thus, the variance and the posterior spread have the same upper bound, which means that the full posterior converges at the same rate as the posterior mean.
@@ C.2.2. Bounding the bias
Let Assumptions 4.2, 4.3, and 4.4 hold. Let the bias $B_{N}$ be as in (68a). There exists $c_{0}>8$, $c\in(0,1/4)$, and $N_{0}\geq 1$ (all independent of $N$ and $f^{\dagger}$) such that for any $N\geq N_{0}$, it holds that
@@
Recalling $\widehat{M}$ from (77) and the identity (109) from Lemma C.21 gives
@@
Thus, the right hand side of the display (78) in Lemma C.8 is bounded above by
$\begin{split}&2\mu^{2}\sum_{k=1}^{\infty}\frac{\sigma_{k}^{\prime}\lambda_{k}}{\sigma_{k}\lambda_{k}+\mu}\Bigg{\lvert}\sum_{j=1}^{\infty}{\frac{\lvert f^{\dagger}_{j}\rvert\lambda_{j}^{-1/2}}{(\sigma_{j}\lambda_{j}+\mu)^{1/2}}\big{\lvert}\langle\varphi_{k},\varphi_{j}\rangle\big{\rvert}}\Bigg{\rvert}^{2}\\ &\quad+2\mu^{2}\sum_{k=1}^{\infty}\frac{\sigma_{k}^{\prime}\lambda_{k}}{\sigma_{k}\lambda_{k}+\mu}\Bigg{\lvert}\sum_{j=1}^{\infty}{\frac{\lvert f^{\dagger}_{j}\rvert\lambda_{j}^{-1/2}}{(\sigma_{j}\lambda_{j}+\mu)^{1/2}}\big{\lvert}\langle\varphi_{k},\widehat{M}\widehat{T}\varphi_{j}\rangle\big{\rvert}}\Bigg{\rvert}^{2}\\ &\quad\quad=\underbrace{2\mu^{2}\sum_{k=1}^{\infty}\frac{\sigma_{k}^{\prime}\lvert f^{\dagger}_{k}\rvert^{2}}{(\sigma_{k}\lambda_{k}+\mu)^{2}}}_{\eqqcolon\overline{B}_{N}}+\underbrace{2\mu^{2}\sum_{k=1}^{\infty}\frac{\sigma_{k}^{\prime}\lambda_{k}}{\sigma_{k}\lambda_{k}+\mu}\Bigg{\lvert}\sum_{j=1}^{\infty}{\frac{\lvert f^{\dagger}_{j}\rvert\lambda_{j}^{-1/2}}{(\sigma_{j}\lambda_{j}+\mu)^{1/2}}\big{\lvert}\langle\varphi_{k},\widehat{M}\widehat{T}\varphi_{j}\rangle\big{\rvert}}\Bigg{\rvert}^{2}}_{\eqqcolon\widehat{B}_{N}}.\end{split}$ (79)
@@
To this end, let $E$ be the event from (107):
$E=\Bigl{\{}\big{\lVert}\widehat{T}\big{\rVert}_{\mathcal{L}(H)}\leq 1/2\Bigr{\}}\,.$ (80)
@@
$E_{0}\coloneqq\bigl{\{}\widehat{B}_{N}\leq\varepsilon\bigr{\}}\,.$ (81)
Let the intersection $E_{0}\cap E$ be our “good” event. On this event, our variance and bias bounds will hold simultaneously. For our results to be meaningful, we must show that $E_{0}\cap E$ has high probability. Since $\operatorname{\mathbb{P}}(E_{0}\,|\,E)=1-\operatorname{\mathbb{P}}(E_{0}^{\textsf{c}}\,|\,E)$, we have
$\begin{split}\operatorname{\mathbb{P}}(E_{0}\cap E)&=\operatorname{\mathbb{P}}(E_{0}\,|\,E)\operatorname{\mathbb{P}}(E)\\ &=\operatorname{\mathbb{P}}(E)-\operatorname{\mathbb{P}}(E_{0}^{\textsf{c}}\,|\,E)\operatorname{\mathbb{P}}(E)\\ &=\operatorname{\mathbb{P}}(E)-\operatorname{\mathbb{P}}(E_{0}^{\textsf{c}}\cap E)\,.\end{split}$ (82)
Thus, to lower bound the probability of $E_{0}\cap E$, it suffices to upper bound
$\operatorname{\mathbb{P}}(E_{0}^{\textsf{c}}\cap E)=\operatorname{\mathbb{P}}\bigl{(}\{\widehat{B}_{N}>\varepsilon\}\cap E\bigr{)}\,.$
On the event $E$, it holds that $\lVert\widehat{M}\rVert_{\mathcal{L}(H)}\leq 2$ by (74). This, the symmetry of $\widehat{M}$, and the Cauchy–Schwarz inequality imply that
@@
on the event $E$. Notice that in the last line of the preceding display, the factor $(\mathcal{I}_{N})^{2}$ is multiplying our high probability upper bound (69) on $V_{N}$. Hence, for the contribution from $\widehat{B}_{N}$ to be negligible relative to the upper bound on $V_{N}$, it suffices to show that $\mathcal{I}_{N}\lesssim 1$ for all sufficiently large $N$ with high probability. Indeed, for some $\tau>0$ to be determined later, choose
@@
in the definition (81) of $E_{0}$. Then the monotonicity of probability measure (i.e., if $A_{1}\subseteq A_{2}$, then $\operatorname{\mathbb{P}}(A_{1})\leq\operatorname{\mathbb{P}}(A_{2})$) implies that
$\operatorname{\mathbb{P}}(E_{0}^{\textsf{c}}\cap E)\leq\operatorname{\mathbb{P}}(\{\mathcal{I}_{N}>\tau\}\cap E)\leq\operatorname{\mathbb{P}}\{\mathcal{I}_{N}>\tau\}\,.$ (84)
@@
That is, $\widehat{T}\varphi_{j}$ is a sum of iid random vectors in the Hilbert space $H$. To show that the scalar random variables $\lVert\widehat{T}\varphi_{j}\rVert$ are subexponential, we first need to control the subexponential norm of the $\lVert\zeta_{jn}\rVert$. The next lemma accomplishes this task.
@@
The second line is due to Jensen’s inequality. Let $u$ be an iid copy of $u_{1}\sim\nu$, so that $v\coloneqq\mathcal{C}_{\mu}^{-1/2}\Lambda^{1/2}u$ is an iid copy of $v_{1}$. By Assumption 4.3 and the assumption (A1) that $\Lambda$ and $\Sigma$ share the orthonormal eigenbasis $\{\varphi_{j}\}$, it holds that
@@
The triangle inequality in the Banach space $L^{\ell/2}_{\operatorname{\mathbb{P}}}(D;\mathbb{R})$ (since $\ell\geq 2$) and the Cauchy–Schwarz inequality yields
@@
We need the following proposition that relies heavily on [59, Theorem 1, p. 144].
@@
holds for some $\sigma>0$ and $b>0$, then the partial sums $\mathcal{S}_{N}\coloneqq\sum_{n=1}^{N}X_{n}$ satisfy the subexponential condition $\lVert\mathcal{S}_{N}\rVert\in\operatorname{SE}(2\sigma^{2},2b)$, that is,
@@
Since $\mathcal{X}$ is a separable Banach space, [59, Theorem 1, p. 144] shows that
@@
But since $\lVert\mathcal{S}_{N}\rVert\geq 0$ a.s., the one-sided Bernstein moment generating function bound [75, Proposition 2.14, Equation (2.22a), p. 31] applied to $-\lVert\mathcal{S}_{N}\rVert$ yields
@@
$\big{\lVert}\widehat{T}\varphi_{j}\big{\rVert}\in\operatorname{SE}\bigg{(}\frac{2\varsigma_{j}^{2}}{N},\frac{2\varsigma_{j}}{N}\bigg{)}\,.$ (92)
@@
Under Assumption 4.2 and 4.4, it holds that
@@
by the Cauchy–Schwarz inequality and Assumption 4.3. The rightmost series converges (because $12(\alpha+s)>2>$ 1 by the last assertion of Assumption 4.4) and is bounded above by a constant (independent of $N$) times
@@
Combining the previous results with those of [16, Appendix B, pp. 30–31] establishes that the entire series $\mathcal{I}_{N}$ (85) is a real-valued subexponential random variable.
@@
Let Assumptions 4.2, 4.3, and 4.4 be satisfied and $\mathcal{I}_{N}$ be defined as in (85). It holds that
$\mathcal{I}_{N}\in\operatorname{SE}\bigg{(}\frac{2}{N}\big{\lVert}w^{(N)}\big{\rVert}_{\ell^{1}(\mathbb{N})}^{2},\frac{2}{N}\big{\lVert}w^{(N)}\big{\rVert}_{\ell^{1}(\mathbb{N})}\bigg{)}\quad\text{for all}\quad N\in\mathbb{N}\,.$ (95)
@@
Fix $N\in\mathbb{N}$. A change of variables in Definition C.2 of subexponential and Lemma C.11 imply that $s_{j}\lVert\widehat{T}\varphi_{j}\rVert\in\operatorname{SE}(2s_{j}^{2}\varsigma_{j}^{2}/N,2s_{j}\varsigma_{j}/N)$ for any $j$. For any $J\in\mathbb{N}$, let $\mathcal{I}_{N}^{(J)}\coloneqq\sum_{j=1}^{J}s_{j}\lVert\widehat{T}\varphi_{j}\rVert$. Even though the summands $\{s_{j}\lVert\widehat{T}\varphi_{j}\rVert\}_{j\in\mathbb{N}}$ are a dependent sequence of random variables, [16, Lemma B.5, p. 30] shows that
$\mathcal{I}_{N}^{(J)}\in\operatorname{SE}\Bigg{(}\frac{2}{N}\bigg{\lvert}\sum_{j=1}^{J}s_{j}\varsigma_{j}\bigg{\rvert}^{2},\frac{2}{N}\sum_{j=1}^{J}s_{j}\varsigma_{j}\Bigg{)}\,.$
@@
Since the right hand side of (96) is finite by Lemma C.12, a monotone convergence argument [16, Lemma B.3, p. 30] shows that
@@
Let Assumptions 4.2, 4.3, and 4.4 be satisfied and $\mathcal{I}_{N}$ be defined as in (85). For any $N\in\mathbb{N}$, it holds that
@@
By Lemma C.13 and [75, Proposition 2.9, p. 26], it holds that
@@
Let Assumptions 4.2, 4.3, and 4.4 be satisfied and $\mathcal{I}_{N}$ be defined as in (85). There exists $c_{0}>1$ and $c\in(0,1/4)$, both independent of $N$ and $f^{\dagger}$, such that
@@
for all $N\in\mathbb{N}$. The right hand side of the preceding display is thus an upper bound to $\operatorname{\mathbb{P}}\{\mathcal{I}_{N}\geq\lVert w^{(N)}\rVert_{\ell^{1}}/\sqrt{2N}+\lVert f^{\dagger}\rVert_{\mathcal{H}^{s}}\}$. To complete the proof, notice that $\lVert w^{(N)}\rVert_{\ell^{1}(\mathbb{N})}/\sqrt{2N}\leq(c^{\prime}/2)\lVert f^{\dagger}\rVert_{\mathcal{H}^{s}}N^{-\min(1,(\alpha+s-1)/(\alpha+p))/2}\leq(c^{\prime}/2)\lVert f^{\dagger}\rVert_{\mathcal{H}^{s}}$ because $\alpha+s>1$ by Assumption 4.4. By monotonicity of probability measure, this implies the asserted result (98) with $c_{0}\coloneqq 1+c^{\prime}/2$. ∎
@@
Recalling the event $E_{0}$ from (81) with $\varepsilon$ as in (83), choose $\tau=c^{\prime}\lVert f^{\dagger}\rVert_{\mathcal{H}^{s}}$ with $c^{\prime}>1$ equal to the constant $c_{0}$ in the hypotheses of Lemma C.15. On the good event $E_{0}\cap E$ from (81) and (80), it holds by (79) that $B_{N}\leq\overline{B}_{N}+\varepsilon$ which is precisely the claimed upper bound (76) with $c_{0}\coloneqq 8(c^{\prime})^{2}$. It remains to lower bound the probability of $E_{0}\cap E$. By (82) and (84), it holds that
$\operatorname{\mathbb{P}}(E_{0}\cap E)=\operatorname{\mathbb{P}}(E)-\operatorname{\mathbb{P}}(E_{0}^{\textsf{c}}\cap E)\geq\operatorname{\mathbb{P}}(E)-\operatorname{\mathbb{P}}(\mathcal{I}_{N}>c^{\prime}\lVert f^{\dagger}\rVert_{\mathcal{H}^{s}})\,.$
By hypothesis, the assertion of Lemma C.20 (i.e., that $\operatorname{\mathbb{P}}(E)\geq 1-e^{-c_{1}N}$ for some $c_{1}\in(0,1)$) holds true provided that $N\geq N_{0}$ with $N_{0}\geq 1$ defined in (106). Combining this with Lemma C.15 shows that, for all $N\geq N_{0}$, the good set satisfies
$\operatorname{\mathbb{P}}(E_{0}\cap E)$ $\geq 1-e^{-c_{1}N}-\exp\Bigl{(}-c_{2}N^{\min\bigl{(}1,\frac{\alpha+s-1}{\alpha+p}\bigr{)}}\Bigr{)}$
@@
for some $c_{2}\in(0,1/4)$. We defined $c\coloneqq\min(c_{1},c_{2})\in(0,1/4)$ in the last line of the preceding display. To complete the proof, notice that if $E_{0}\cap E$ occurs, then $E$ also occurs. But the variance bound (69) also holds true on $E$ (as shown in the proof of Proposition C.5). This proves the final assertion of the proposition. ∎

### v2

@@ C.2. Proofs for end-to-end learning of general linear functionals
$B_{N}$ $\coloneqq\big{\lVert}(\Sigma^{\prime})^{1/2}(\operatorname{Id}_{H}-A_{N}S_{N})f^{\dagger}\big{\rVert}^{2}\quad\text{and}\quad$ (73a)
@@ C.2.1. Bounding the variance
The expression for the variance upper bound in Proposition C.5 is in precisely the same form as that found in [17, Equation (A.1b), p. 24]. To prove the proposition, we require some preparatory results. First, notice that
$\operatorname{\mathbb{E}}\mathopen{}\mathclose{{}\left[\big{\lVert}(\Sigma^{\prime})^{1/2}A_{N}\Xi\big{\rVert}^{2}\,\big{|}\,U}\right]=\operatorname{\mathbb{E}}^{g\sim\mathcal{N}(0,\operatorname{Id}_{{\mathbb{R}^{N}}})}\big{\lVert}(\Sigma^{\prime})^{1/2}A_{N}g\big{\rVert}^{2}=\operatorname{tr}\big{(}(\Sigma^{\prime})^{1/2}A_{N}^{\phantom{*}}A_{N}^{*}(\Sigma^{\prime})^{1/2}\big{)}\,.$
Now using the fact that $\widehat{\Sigma}=S_{N}^{*}S_{N}^{\phantom{*}}/N$ from (14) and defining
@@
$A_{N}^{\phantom{*}}A_{N}^{*}=\frac{1}{N}\biggl{[}\Lambda^{1/2}\biggl{(}\widehat{\mathcal{C}}+\frac{\gamma^{2}}{N}\operatorname{Id}_{H}\biggr{)}^{-1}\widehat{\mathcal{C}}\biggl{(}\widehat{\mathcal{C}}+\frac{\gamma^{2}}{N}\operatorname{Id}_{H}\biggr{)}^{-1}\Lambda^{1/2}\biggr{]}\,.$
@@
The proof of Proposition C.5 also justifies the claim made in Remark 4.10. Indeed, the second term on the right-hand side of the equality (37) is the posterior spread with respect to the weighted norm (18). Equation (78) upper bounds the posterior spread, which in turn upper bounds the variance (73b) in the bias–variance decomposition of (18). Thus, the variance and the posterior spread have the same upper bound.
@@ C.2.2. Bounding the bias
Let Assumptions 4.2, 4.3, and 4.5 hold. Let the bias $B_{N}$ be as in (73a). There exists $c_{0}>8$, $c\in(0,1/4)$, and $N_{0}\geq 1$ (all independent of $N$ and $f^{\dagger}$) such that for any $N\geq N_{0}$, it holds that
@@
Recalling $\widehat{M}$ from (82) and using the identity (115) from Lemma C.22 gives
@@
Thus, the right-hand side of the display (83) in Lemma C.8 is bounded above by
$\begin{split}&2\mu^{2}\sum_{k=1}^{\infty}\frac{\sigma_{k}^{\prime}\lambda_{k}}{\sigma_{k}\lambda_{k}+\mu}\Bigg{\lvert}\sum_{j=1}^{\infty}{\frac{\lvert f^{\dagger}_{j}\rvert\lambda_{j}^{-1/2}}{(\sigma_{j}\lambda_{j}+\mu)^{1/2}}\big{\lvert}\langle\varphi_{k},\varphi_{j}\rangle\big{\rvert}}\Bigg{\rvert}^{2}\\ &\qquad\qquad+2\mu^{2}\sum_{k=1}^{\infty}\frac{\sigma_{k}^{\prime}\lambda_{k}}{\sigma_{k}\lambda_{k}+\mu}\Bigg{\lvert}\sum_{j=1}^{\infty}{\frac{\lvert f^{\dagger}_{j}\rvert\lambda_{j}^{-1/2}}{(\sigma_{j}\lambda_{j}+\mu)^{1/2}}\big{\lvert}\langle\varphi_{k},\widehat{M}\widehat{T}\varphi_{j}\rangle\big{\rvert}}\Bigg{\rvert}^{2}\\ &=\underbrace{2\mu^{2}\sum_{k=1}^{\infty}\frac{\sigma_{k}^{\prime}\lvert f^{\dagger}_{k}\rvert^{2}}{(\sigma_{k}\lambda_{k}+\mu)^{2}}}_{\eqqcolon\overline{B}_{N}}+\underbrace{2\mu^{2}\sum_{k=1}^{\infty}\frac{\sigma_{k}^{\prime}\lambda_{k}}{\sigma_{k}\lambda_{k}+\mu}\Bigg{\lvert}\sum_{j=1}^{\infty}{\frac{\lvert f^{\dagger}_{j}\rvert\lambda_{j}^{-1/2}}{(\sigma_{j}\lambda_{j}+\mu)^{1/2}}\big{\lvert}\langle\varphi_{k},\widehat{M}\widehat{T}\varphi_{j}\rangle\big{\rvert}}\Bigg{\rvert}^{2}}_{\eqqcolon\widehat{B}_{N}}.\end{split}$ (84)
@@
To this end, let $\mathsf{E}$ be the event from (112):
$\mathsf{E}=\Bigl{\{}\big{\lVert}\widehat{T}\big{\rVert}_{\mathcal{L}(H)}\leq 1/2\Bigr{\}}\,.$ (85)
@@
$\mathsf{E}_{0}\coloneqq\bigl{\{}\widehat{B}_{N}\leq\varepsilon\bigr{\}}\,.$ (86)
Let the intersection $\mathsf{E}_{0}\cap\mathsf{E}$ be our “good” event. On this event, our variance and bias bounds will hold simultaneously. For our results to be meaningful, we must show that $\mathsf{E}_{0}\cap\mathsf{E}$ has high probability. Since $\operatorname{\mathbb{P}}(\mathsf{E}_{0}\,|\,\mathsf{E})=1-\operatorname{\mathbb{P}}(\mathsf{E}_{0}^{\textsf{c}}\,|\,\mathsf{E})$, we have
$\begin{split}\operatorname{\mathbb{P}}(\mathsf{E}_{0}\cap\mathsf{E})&=\operatorname{\mathbb{P}}(\mathsf{E}_{0}\,|\,\mathsf{E})\operatorname{\mathbb{P}}(\mathsf{E})\\ &=\operatorname{\mathbb{P}}(\mathsf{E})-\operatorname{\mathbb{P}}(\mathsf{E}_{0}^{\textsf{c}}\,|\,\mathsf{E})\operatorname{\mathbb{P}}(\mathsf{E})\\ &=\operatorname{\mathbb{P}}(\mathsf{E})-\operatorname{\mathbb{P}}(\mathsf{E}_{0}^{\textsf{c}}\cap\mathsf{E})\,.\end{split}$ (87)
Thus, to lower bound the probability of $\mathsf{E}_{0}\cap\mathsf{E}$, it suffices to upper bound
$\operatorname{\mathbb{P}}(\mathsf{E}_{0}^{\textsf{c}}\cap\mathsf{E})=\operatorname{\mathbb{P}}\bigl{(}\{\widehat{B}_{N}>\varepsilon\}\cap\mathsf{E}\bigr{)}\,.$
On the event $\mathsf{E}$, it holds that $\lVert\widehat{M}\rVert_{\mathcal{L}(H)}\leq 2$ by (79). This, the symmetry of $\widehat{M}$, and the Cauchy–Schwarz inequality imply that
@@
on the event $\mathsf{E}$. Notice that in the last line of the preceding display, the factor $(\mathcal{I}_{N})^{2}$ is multiplying our high probability upper bound (74) on $V_{N}$. Hence, for the contribution from $\widehat{B}_{N}$ to be negligible relative to the upper bound on $V_{N}$, it suffices to show that $\mathcal{I}_{N}\lesssim 1$ for all sufficiently large $N$ with high probability. Indeed, for some $\tau>0$ to be determined later, choose
@@
in the definition (86) of $\mathsf{E}_{0}$. Then the monotonicity of probability measure (i.e., if $\mathsf{A}_{1}\subseteq\mathsf{A}_{2}$, then $\operatorname{\mathbb{P}}(\mathsf{A}_{1})\leq\operatorname{\mathbb{P}}(\mathsf{A}_{2})$) implies that
$\operatorname{\mathbb{P}}(\mathsf{E}_{0}^{\textsf{c}}\cap\mathsf{E})\leq\operatorname{\mathbb{P}}(\{\mathcal{I}_{N}>\tau\}\cap\mathsf{E})\leq\operatorname{\mathbb{P}}\{\mathcal{I}_{N}>\tau\}\,.$ (89)
@@
That is, $\widehat{T}\varphi_{j}$ is a sum of i.i.d. random vectors in the Hilbert space $H$. To show that the scalar random variables $\lVert\widehat{T}\varphi_{j}\rVert$ are subexponential, we first need to control the subexponential norm of the $\lVert\zeta_{jn}\rVert$. The next lemma accomplishes this task.
@@
The second line is due to Jensen’s inequality. Let $u$ be an i.i.d. copy of $u_{1}\sim\nu$, so that $v\coloneqq\mathcal{C}_{\mu}^{-1/2}\Lambda^{1/2}u$ is an i.i.d. copy of $v_{1}$. By Assumption 4.3 and the assumption (A1) that $\Lambda$ and $\Sigma$ share the orthonormal eigenbasis $\{\varphi_{j}\}$, it holds that
@@
The triangle inequality in the Banach space $L^{\ell/2}_{\operatorname{\mathbb{P}}}(\Omega;\mathbb{R})$ (since $\ell\geq 2$) and the Cauchy–Schwarz inequality yields
@@
We need the following proposition that relies heavily on [70, Theorem 1, p. 144].
@@
holds for some $\sigma>0$ and $b>0$, then the partial sums $\mathcal{S}_{N}\coloneqq\sum_{n=1}^{N}X_{n}$ satisfy the subexponential condition $\lVert\mathcal{S}_{N}\rVert\in\operatorname{\mathsf{SE}}(2\sigma^{2},2b)$, that is,
@@
Since $\mathcal{X}$ is a separable Banach space, [70, Theorem 1, p. 144] shows that
@@
But since $\lVert\mathcal{S}_{N}\rVert\geq 0$ almost surely, the one-sided Bernstein moment generating function bound [87, Proposition 2.14, Equation (2.22a), p. 31] applied to $-\lVert\mathcal{S}_{N}\rVert$ yields
@@
$\big{\lVert}\widehat{T}\varphi_{j}\big{\rVert}\in\operatorname{\mathsf{SE}}\bigg{(}\frac{2\varsigma_{j}^{2}}{N},\frac{2\varsigma_{j}}{N}\bigg{)}\,.$ (97)
@@
Under Assumption 4.2 and 4.5, it holds that
@@
by the Cauchy–Schwarz inequality. The rightmost series converges (because $12(\alpha+s)>2>$ 1 by the last assertion of Assumption 4.5) and is bounded above by a constant (independent of $N$) times
@@
Combining the previous results with those of [17, Appendix B, pp. 30–31] establishes that the entire series $\mathcal{I}_{N}$ (90) is a real-valued subexponential random variable.
@@
Let Assumptions 4.2, 4.3, and 4.5 be satisfied and $\mathcal{I}_{N}$ be defined as in (90). It holds that
$\mathcal{I}_{N}\in\operatorname{\mathsf{SE}}\bigg{(}\frac{2}{N}\big{\lVert}w^{(N)}\big{\rVert}_{\ell^{1}(\mathbb{N})}^{2},\frac{2}{N}\big{\lVert}w^{(N)}\big{\rVert}_{\ell^{1}(\mathbb{N})}\bigg{)}\quad\text{for all}\quad N\in\mathbb{N}\,.$ (100)
@@
Fix $N\in\mathbb{N}$. A change of variables in Definition C.2 of subexponential and Lemma C.11 imply that $s_{j}\lVert\widehat{T}\varphi_{j}\rVert\in\operatorname{\mathsf{SE}}(2s_{j}^{2}\varsigma_{j}^{2}/N,2s_{j}\varsigma_{j}/N)$ for any $j$. For any $J\in\mathbb{N}$, let $\mathcal{I}_{N}^{(J)}\coloneqq\sum_{j=1}^{J}s_{j}\lVert\widehat{T}\varphi_{j}\rVert$. Even though the summands $\{s_{j}\lVert\widehat{T}\varphi_{j}\rVert\}_{j\in\mathbb{N}}$ are a dependent sequence of random variables, [17, Lemma B.5, p. 30] shows that
$\mathcal{I}_{N}^{(J)}\in\operatorname{\mathsf{SE}}\Bigg{(}\frac{2}{N}\bigg{\lvert}\sum_{j=1}^{J}s_{j}\varsigma_{j}\bigg{\rvert}^{2},\frac{2}{N}\sum_{j=1}^{J}s_{j}\varsigma_{j}\Bigg{)}\,.$
@@
Since the right-hand side of (101) is finite by Lemma C.12, a monotone convergence argument [17, Lemma B.3, p. 30] shows that
@@
Let Assumptions 4.2, 4.3, and 4.5 be satisfied and $\mathcal{I}_{N}$ be defined as in (90). For any $N\in\mathbb{N}$, it holds that
@@
By Lemma C.13 and [87, Proposition 2.9, p. 26], it holds that
@@
Let Assumptions 4.2, 4.3, and 4.5 be satisfied and $\mathcal{I}_{N}$ be defined as in (90). There exists $c_{0}>1$ and $c\in(0,1/4)$, both independent of $N$ and $f^{\dagger}$, such that
@@
for all $N\in\mathbb{N}$. The right-hand side of the preceding display is thus an upper bound to $\operatorname{\mathbb{P}}\{\mathcal{I}_{N}\geq\lVert w^{(N)}\rVert_{\ell^{1}}/\sqrt{2N}+\lVert f^{\dagger}\rVert_{\mathcal{H}^{s}}\}$. To complete the proof, notice that $\lVert w^{(N)}\rVert_{\ell^{1}(\mathbb{N})}/\sqrt{2N}\leq(c^{\prime}/2)\lVert f^{\dagger}\rVert_{\mathcal{H}^{s}}N^{-\min(1,(\alpha+s-1)/(\alpha+p))/2}\leq(c^{\prime}/2)\lVert f^{\dagger}\rVert_{\mathcal{H}^{s}}$ because $\alpha+s>1$ by Assumption 4.5. By monotonicity of probability measure, this implies the asserted result (103) with $c_{0}\coloneqq 1+c^{\prime}/2$. ∎
@@
Recalling the event $\mathsf{E}_{0}$ from (86) with $\varepsilon$ as in (88), choose $\tau=c^{\prime}\lVert f^{\dagger}\rVert_{\mathcal{H}^{s}}$ with $c^{\prime}>1$ equal to the constant $c_{0}$ in the hypotheses of Lemma C.15. On the good event $\mathsf{E}_{0}\cap\mathsf{E}$ from (86) and (85), it holds by (84) that $B_{N}\leq\overline{B}_{N}+\varepsilon$ which is precisely the claimed upper bound (81) with $c_{0}\coloneqq 8(c^{\prime})^{2}$. It remains to lower bound the probability of $\mathsf{E}_{0}\cap\mathsf{E}$. By (87) and (89), it holds that
$\operatorname{\mathbb{P}}(\mathsf{E}_{0}\cap\mathsf{E})=\operatorname{\mathbb{P}}(\mathsf{E})-\operatorname{\mathbb{P}}(\mathsf{E}_{0}^{\textsf{c}}\cap\mathsf{E})\geq\operatorname{\mathbb{P}}(\mathsf{E})-\operatorname{\mathbb{P}}(\mathcal{I}_{N}>c^{\prime}\lVert f^{\dagger}\rVert_{\mathcal{H}^{s}})\,.$
By hypothesis, the assertion of Lemma C.20 (i.e., that $\operatorname{\mathbb{P}}(\mathsf{E})\geq 1-e^{-c_{1}N}$ for some $c_{1}\in(0,1)$) holds true provided that $N\geq N_{0}$ with $N_{0}\geq 1$ defined in (111). Combining this with Lemma C.15 shows that, for all $N\geq N_{0}$, the good set satisfies
$\operatorname{\mathbb{P}}(\mathsf{E}_{0}\cap\mathsf{E})$ $\geq 1-e^{-c_{1}N}-\exp\Bigl{(}-c_{2}N^{\min\bigl{(}1,\frac{\alpha+s-1}{\alpha+p}\bigr{)}}\Bigr{)}$
@@
for some $c_{2}\in(0,1/4)$. We defined $c\coloneqq\min(c_{1},c_{2})\in(0,1/4)$ in the last line of the preceding display. To complete the proof, notice that if $\mathsf{E}_{0}\cap\mathsf{E}$ occurs, then $\mathsf{E}$ also occurs. But the variance bound (74) also holds true on $\mathsf{E}$ (as shown in the proof of Proposition C.5). This proves the final assertion of the proposition. ∎